OpenStudy (anonymous):

4x^2*squareroot(8x^3

OpenStudy (shamil98):

\[4x^2\sqrt{8x^3}\] is that your problem?

OpenStudy (anonymous):

yes sir

OpenStudy (shamil98):

First, let's start off by simplifying the √8x^3 what is the square root of √8 ?

OpenStudy (anonymous):

2.83

OpenStudy (shamil98):

Write it in radicals, √8 = √4 x √2 =

OpenStudy (anonymous):

2x squareroot2x

OpenStudy (anonymous):

x^3 squreroot = x,x

OpenStudy (shamil98):

yes so now we have 4x^2(2√2x^3)

OpenStudy (shamil98):

Simplify the like terms 8x^2(√2x^3)

OpenStudy (anonymous):

i have 4x^2*2x squareroot 2x

OpenStudy (shamil98):

\[4x^2 2\sqrt{2x^3}\] \[= 8x^2\sqrt{2x^3}\]

OpenStudy (anonymous):

but cant you take out x^2 out of the x^3

OpenStudy (shamil98):

I'm getting to that part , now we simplify the √x^3

OpenStudy (anonymous):

x squareroot x

OpenStudy (shamil98):

\[\sqrt[]{x^3} = \sqrt{x^2} \times x\]

OpenStudy (anonymous):

so we get